OpenStudy (maeganr):

APPLICATIONS OF LINEAR SYSTEMS: 3 Word Problems! Please Help Asap! Due tomorrow!

OpenStudy (maeganr):

#4. One plane at 520 feet is ascending at a rate of 40 feet per minute, while another plane at 3800 feet is defending at a rate of 120 feet per minute. How long will it take the two planes to be at the same altitude? #6. You are getting ready to move and have some friends to help. For lunch, you buy the following sandwiches at the local deli for $30: six tuna sandwiches and six turkey sandwiches. Later at night, everyone is hungry again and you buy four tuna sandwiches and eight turkey sandwiches for $30.60. What is the price of each sandwich? #8. At an all-you-can-eat barbecue fundraiser that you are sponsoring, adults pay $6 for a dinner and children pay $4 for a dinner. 212 people attend and you raise $1128. What is the total number of adults and the total number of children attending? a) What is the system of equations that you can use to solve this problem? b) What method would you use to solve the system? Why?

OpenStudy (maeganr):

@Butterfly16 ...

OpenStudy (anonymous):

this sounds like easy kinematics xD For plane one \[y_{1-0}=520ft\] \[v_{1-0}=40 \dfrac{ft} {min}\] For plane two \[y_{2-0}=3800ft\] \[v_{2-0}=-120 \dfrac{ft} {min}\] When the two planes are at the same height y1(t)=y2(t), so you equal their postioning equations \[y(t)=y_{0}+v_{0}(t_{2}-t{1})+\dfrac{1} {2} a(t_{2}-t_{1})^2\] We don't have acceleration and t1=0 so, we will have \[t_{2}=t\] And our equations will be in the form \[y_{t}=y_{0}+v_{0}t\] Now, as we want plane one and plane two to be at the same height, They'll have the same "y(t)" and we can equal their equations \[520ft+40 \dfrac{ft} {min}t=3800ft-120 \dfrac{ft} {min}t\] \[t=20.5 min\]

OpenStudy (maeganr):

Whats "kinematics*"?^^^

OpenStudy (maeganr):

And what? Is that? .......

OpenStudy (anonymous):

For #6 you write a system, x= tuna sandwich and y=turkey sandwiches 6x+6y=30 4x+8y=30.6 So x=2.35 and y=2.65 As for your questions, kinematics is physics simple stuff (movement), ft=feet, min=minute, the position equations are also called kinematic equations, they describe the positioning of a body "against" time, and are actually vectorial equations... y(t) is the position of your body in time t

OpenStudy (maeganr):

Oh. I knew that.....

OpenStudy (anonymous):

So, you got it?

OpenStudy (maeganr):

Yeah, I think. And really, thanks for the help like everryyyyyyy day. (:

OpenStudy (anonymous):

And for the last one x= number of adults y=number of children a)4x+6y=1128 x+y=212 b)I dunno, i'll use substitution x=72 and y=140